### v1

# Do Comments and Expertise Still Matter? An Experiment on Programmers’ Adoption of AI-Generated JavaScript Code

Abstract
This paper investigates the factors influencing programmers’ adoption of AI-generated JavaScript code recommendations. It extends prior research by (1) utilizing objective (as opposed to the typically self-reported) measurements for programmers’ adoption of AI-generated code and (2) examining whether AI-generated comments added to code recommendations and development expertise drive AI-generated code adoption. We tested these potential drivers in an online experiment with 173 programmers. Participants were asked to answer some questions to demonstrate their level of development expertise. Then, they were asked to solve a LeetCode problem without AI support. After attempting to solve the problem on their own, they received an AI-generated solution to assist them in refining their solutions. The solutions provided were manipulated to include or exclude AI-generated comments (a between-subjects factor). Programmers’ adoption of AI-generated code was gauged by code similarity between AI-generated solutions and participants’ submitted solutions, providing a more reliable and objective measurement of code adoption behaviors. Our findings revealed that the presence of comments significantly influences programmers’ adoption of AI-generated code regardless of the participants’ development expertise.
keywords:
AI programming assistant , empirical software engineering , technology adoption , human-computer interaction
journal: Journal of Systems and Software
[…]
[label1]organization=The University of Melbourne, addressline=Grattan Street, city=Melbourne, postcode=3010, state=Victoria, country=Australia
[…]
[label2]organization=Singapore Management University, addressline=81 Victoria Street, city=Singapore, postcode=188065, state=Singapore, country=Singapore
1 Introduction
One of the most common hurdles in software engineering is turning ideas into programs [1]. A code generation system capable of automatically translating natural languages into code could assist programmers, leading to a more efficient and productive programming paradigm [2, 3]. Automatically generating code from high-level descriptions has been a historic challenge for decades in the computer science community [4]. The recent breakthrough in artificial intelligence, especially in large language models (LLMs), brings the computer science community closer to addressing this challenge [2, 5]. AI has promising potential to revolutionize software engineering processes, including code search, code recommendation, automatic bug fixing, code review, risk assessment of code changes, and troubleshooting [6].
Despite the fact that recent advancements in artificial intelligence are reshaping software engineering [5, 7], little research has investigated the factors that influence programmers’ adoption of AI-generated code [8]. Ge & Wu [9] conducted an interview-based qualitative study to investigate the factors that impact the adoption of ChatGPT for bug fixing. They discovered several key influential factors, including performance expectancy, effort expectancy, social influence, facilitating conditions, data security, and trust [9]. Russo conducted a survey-based study to investigate the factors influencing developers’ adoption of generative AI tools in software engineering [8]. Russo proposed the Human-AI Collaboration and Adaptation Framework (HACAF), a theoretical framework to explain programmers’ adoption of AI coding tools [8]. Liang et al. conducted a large-scale survey to evaluate the reasons programmers choose to use or avoid AI coding assistants by directly asking participants to answer why they use or refuse AI coding tools [10].
While previous studies exploring the factors influencing programmers’ adoption of AI-generated code have made valuable contributions, several research gaps remain. The first gap pertains to the over-reliance on self-reported data. The assessment of participants’ adoption of AI was solely based on self-reported data, lacking complementary objective measurements. Previous studies used participants’ self-reported intentions to gauge their adoption of AI-generated code or AI coding tools. For example, Russo used the question, ”How likely are you to use a language model in your work in the next six months?” to measure participants’ willingness to adopt AI coding tools [8]. The reasons participants adopt or avoid AI-generated code were also collected through self-reported data by directly querying the reasons rather than conducting controlled experiments [10, 8]. Some researchers [11, 12, 13] have raised concerns about the limitations of self-reported questions in surveys. Self-reported data may not fully align with the actual usage pattern since self-reported data are subjective perceptions rather than objective facts.
The second research gap is the non-comprehensive scope of the evaluated factors. Previous studies have primarily examined how productivity and programmers’ perceptions influence the adoption of AI-generated code, neglecting some potentially crucial characteristics of programmers and the generated code. For instance, factors such as the presence of comments in AI-generated code and programmers’ development expertise have been ignored.
To address the first research gap, we have designed a survey-based experiment employing objective measurements to assess participants’ adoption of AI coding suggestions. In the survey, participants were required to complete a JavaScript programming task with the assistance of a code snippet generated by AI. This practical approach allows us to objectively assess their adoption of AI-generated code based on their modifications and usage of the provided snippet. Instead of merely asking participants to express their intentions to adopt AI-generated code, we analyze the submitted code to objectively determine adoption in controlled scenarios. We then conduct statistical analyses, treating potential influential factors as independent variables and the degree of adoption as the dependent variable.
To bridge the second research gap, we undertook a systematic classification of the potential influential factors and evaluated the influence of some unexplored factors. We categorize the potential influential factors into two groups: (A) program factors and (B) programmer factors. Program factors refer to the factors inherent to the program itself, representing the characteristics of the AI-generated code. Programmer factors are tied to the attributes and conditions of the individual programmer. To keep the number of factors manageable in this study, we chose to investigate the presence of comments in the AI-generated code as the program factor and the programmers’ development expertise as the programmer factor.
Our methodology not only deepens our understanding of the determinants of AI code adoption but also ensures the reliability of our findings through more objective measures.
2 Related Work
In this section, we introduce existing studies investigating factors that influence programmers’ adoption of AI-generated code and AI coding tools. We also introduce the independent variables in this research: the presence of comments and development expertise, along with the control variable of attitude toward AI code generators. Since the influence of comments and development expertise on programmers’ adoption of AI-generated code has not been studied before, we focus on introducing related work on their influence on code comprehension.
2.1 Factors Influencing Programmers’ Adoption
There is very little research examining the factors that influence programmers’ adoption of AI code generators or AI-generated code. Previous related studies used interviews and surveys to investigate what impacts participants’ adoption. Ge and Wu conducted semi-structured interviews with 50 participants to scrutinize factors influencing programmers’ adoption of ChatGPT for bug fixing [9]. The Unified Theory of Acceptance and Use of Technology (UTAUT) [14], a model used to explain users’ intentions to adopt a technology, was adopted by Ge and Wu as their theoretical foundation [9]. Their interviews included both open-ended and focused questions about how the four determinants in UTAUT, including performance expectancy, effort expectancy, social influence, and facilitating conditions, affect participants’ adoption of AI-generated bug-fixing suggestions [9]. Their study indicated that all of the four determinants in UTAUT, together with data security and programmers’ trust, could influence the adoption of ChatGPT for bug fixing [9].
Russo conducted a survey-based study with 183 participants, using nine open-ended questions to investigate the factors influencing developers’ adoption of generative AI tools in software engineering [8]. Russo proposed the Human-AI Collaboration and Adaptation Framework (HACAF) to explain the adoption of generative AI tools in software engineering. HACAF is a theoretical model that combines insights from the Technology Acceptance Model (TAM) [15], Diffusion of Innovation (DOI) [16], and Social Cognitive Theory (SCT) [17] to explain programmers’ adoption of generative AI tools. HACAF emphasizes four determinant factors to programmers’ adoption of AI coding tools, including perceptions about the technology, compatibility factors, social factors, and personal and environmental factors [8]. Perceptions about the technology include perceived utility and ease of use. Compatibility factors measure the extent to which AI coding tools are compatible with current software engineering practices. Social factors include social influence and self-efficacy. Personal and environmental factors are personal willingness to experiment with AI coding tools and organizational support.
Liang et al. also explored why programmers use or do not use AI code generators by conducting a survey-based study with 410 participants [10]. Different from Russo’s study [8], Liang et al. employed Likert-scale questions to assess the importance of various determinants that either motivate or discourage participants from using AI coding assistants [10]. Their study indicates that the most significant reasons developers use AI code generators are the capability of AI tools to enhance productivity, including reducing keystrokes, accelerating programming, and assisting with syntax recalling. The pivotal reasons developers refuse to use AI coding tools are the inability of AI-generated code to fulfill specific requirements and the lack of control over producing desired outputs [10].
2.2 Comments
Although the impact of comments on the adoption of AI-generated code has not been systematically studied, related work on the impact of comments on code comprehension inspired our work. The significance of comments is emphasized in many coding guidelines because comments have a profound impact on code comprehension [18]. Several empirical studies have explored the influence of comments on code comprehension [19, 20, 21]. An early empirical study conducted by Takang et al. investigated whether the presence of comments influences code comprehension. Its findings demonstrated that comments play a profound role in enhancing program comprehension, regardless of the identifier naming convention [19]. Nurvitadhi et al. expanded the study by examining how different types of comments influenced code comprehension [20]. Nurvitadhi et al. investigated how class comments and method comments influence Java code understanding. Method comments notably enhance comprehension, while class comments do not have a significant impact [20]. Börstler and Paech further expanded the research by exploring how the quality of comments influences code comprehension [21]. Their finding suggested that the mere presence of comments could enhance programmers’ perception of the readability of the code regardless of the quality of the comments [21].
2.3 Development Expertise
While previous studies have not thoroughly evaluated how development expertise influences programmers’ adoption of AI-generated code, related research on its impact on code comprehension offers a valuable starting point. Development expertise comprises multiple aspects, including general programming experience, programming experience in particular languages, familiarity with task-specific knowledge, and personality traits [22]. Previous studies have explored the influence of development expertise on code comprehension [23, 24]. Lawrie et al. discovered that increased years of work experience and schooling resulted in a better capability of comprehending source code [23]. Wagner and Wyrich studied the influence of intelligence and personality traits on code comprehension [24]. The results of their study showed that there was a strong positive correlation between participants’ general intelligence and their performance on code comprehension and a weaker negative correlation between participants’ conscientiousness and code comprehension [24].
2.4 Attitudes Towards AI Coding Tools
Previous studies show that programmers have multifaceted attitudes toward AI code generators, including positive attitudes [25, 26] and negative attitudes [7, 25]. Several studies demonstrate that some programmers trust AI code generators and believe that AI code generators could improve their productivity [25, 26]. Research has demonstrated that programmers adopt AI coding tools because they trust that AI coding tools could increase productivity [10, 8]. On the other hand, besides positive attitudes, programmers hold negative attitudes as well [7, 25]. Feng et al. examined programmers’ emotional responses to ChatGPT’s code-generation capabilities by gathering social media posts regarding ChatGPT’s code-generation features from Twitter and Reddit [7]. Their analysis showed that the most common sentiments towards ChatGPT were fear, while happiness and anger were the two least common ones [7]. Feng et al. concluded that programmers’ fear derives from two sources, which are the opacity of ChatGPT and the sense of job threat [7].
Kuhail et al. investigated programmers’ fear of job loss [25]. They investigated to what extent programmers fear job loss caused by AI code generators and why they fear or do not fear potential job loss. Participants in their study were required to rate their sense of current and future job security threats. The results suggest that only 24.2% of participants consider AI code generators as non-trivial threats, categorizing them as high threats, moderate threats, or some threats. However, the outlook changes when considering the future, with approximately 50% of the participants perceiving AI code generators as non-trivial threats to their job security. Kuhail et al. also discovered a positive correlation between programmers’ trust in AI code generators and their fear of job replacements [25].
While prior research has investigated various factors influencing programmers’ adoption of AI-generated code and AI coding tools, none have utilized objective measurements for participants’ adoption of AI-generated code. In this work, we focus on how two previously unexplored factors – AI-generated comments and developer expertise – affect the adoption of AI-generated code. By conducting an experiment, we utilized objective measurements to gauge participants’ adoption of AI-generated code and examine how these factors impact programmers’ adoption behaviors, thereby addressing research gaps.
3 Research Methodology
3.1 Research Questions
In this study, we aim to evaluate the influence of two unexplored variables on programmers’ adoption of AI-generated code: the presence of comments in the code (i.e., a program factor) and the programmers’ development expertise (i.e., a programmer factor). We formulate the research questions $\mathbf{RQ_{1}}$ and $\mathbf{RQ_{2}}$ accordingly.
$\mathbf{RQ_{1}}$. How does ($A_{1}$) presence of comments influence programmers’ adoption of AI-generated code?
Previous research has revealed the pivotal role of comments in code comprehension [18]. We hypothesize that comments still matter in programmers’ adoption of AI-generated code. We assume that the presence of comments in AI-generated code would improve programmers’ comprehension of the code, thereby leading to higher rates of code adoption. Our formal hypothesis for RQ1 is:
H1: The factor ($A_{1}$) presence of comments increases programmers’ adoption of AI-generated code.
$\mathbf{RQ_{2}}$. How does ($B_{1}$) development expertise influence programmers’ adoption of AI-generated code?
Programmers with richer development expertise tend to be more experienced in programming and more familiar with task-specific knowledge [22], which could potentially enable them to have a better understanding of AI-generated code without comments. We hypothesize that the influence of ($A_{1}$) presence of comments on programmers’ adoption of AI-generated code would be moderated by ($B_{1}$) development expertise, such that the positive effect of comments on adoption decreases as development expertise increases. The formal hypothesis for RQ2 is:
H2: The positive effect of ($A_{1}$) presence of comments on programmers’ adoption of AI-generated code decreases as a function of ($B_{1}$) development expertise.
3.2 Experimental Design
3.2.1 Experiment Structure
To ensure the reliability and objectivity of our findings, we designed a survey-based experiment with a structured approach. By having participants first solve a programming problem on their own, we create a baseline for their understanding and approach. Then, by introducing an AI-generated solution with or without comments, we can objectively measure how these factors influence their adoption behaviors. This methodology allows us to isolate the effects of comments and development expertise, providing a clear understanding of how these specific factors impact programmers’ willingness to adopt AI-generated code. Additionally, this design discourages participants from passively accepting AI-generated solutions, prompting them to critically evaluate and potentially integrate them with their own coding efforts. Conducting the experiment online with a consistent environment helps maintain control over external variables, ensuring that our results are robust and generalizable.
We developed a custom web application to host the survey on https://survey.changwen-software-engineering.xyz. We conducted pilot tests with five professional developers to ensure that there were no technical errors or usability issues in the survey. Participants in the pilot tests were recruited through the authors’ professional network to ensure that they treated the study seriously and provided quality feedback. We conducted each run with a different participant and fixed the problems encountered in each run.
The survey included six pages in total. The outline of the survey structure is presented in Fig. 1. The first page requested the participants’ consent to participate in the study. Page one provided statements of consent forms, researchers’ contact information, and a link to the plain language statement (PLS). After participants had opted to participate in the survey, the survey would navigate to the second page.
Figure 1: The outline of the survey structure
The second page collected participants’ demographics to obtain the raw data of the (B) programmer factors, including ($B_{11}$) programming experience, ($B_{12}$) programming language proficiency, and ($B_{4}$) attitudes toward AI code generators. One of the programmer factors, ($B_{13}$) domain knowledge familiarity, would be investigated on the fifth page rather than the second page. Domain knowledge is the knowledge utilized to solve the programming problem on the third and fourth pages. We did not present questions about domain knowledge familiarity on page two to avoid potential bias. Asking about domain knowledge before presenting the programming problem might give participants hints, which could affect their responses. For instance, if solving a programming problem required knowledge of the data structure tree, then the domain knowledge familiarity question would ask the participants how familiar they were with the data structure tree. Asking this question would give participants hints that solving the programming problem required using the data structure tree, which could unintentionally make it easier for the participants to understand the provided snippets.
On the third page, we showed participants the programming problems and required them to complete the programming problem in JavaScript on their own. We required participants to solve the programming problem independently first for two reasons. Firstly, it motivated participants to actively engage by trying to solve the problem themselves. In pilot tests, we observed that without this step, many participants would copy AI-generated solutions without fully comprehending the questions. Secondly, it helped us exclude potentially invalid responses, as participants unwilling to put in effort tended to quit the survey at this stage. This self-exclusion helped filter out less meaningful responses and improve data quality. We limited the programming language to JavaScript due to the limited participant number, ensuring manageable feasibility in the study.
The screenshot of the third page is shown in Fig. 2. To provide participants with a coding experience similar to Microsoft Virtual Studio Code (VSCode) [27], we integrated Monaco Editor, a code editor that empowers VSCode [28], into the survey website. Our code editor facilitated syntax highlighting and auto-completion features akin to VSCode, enabling participants to quickly adapt to the coding environment in our survey. Participants could execute their code with test cases specified by themselves or with predefined test cases. There were ten different predefined test cases for each programming problem, and these predefined test cases were hidden from the participants. Participants could view comprehensive execution results, such as error types, error messages, and the line number where an error occurs. A time limit was imposed for solving the programming problem on this page, ranging from a minimum of one minute to a maximum of ten minutes. The lower time limit was set to encourage participants to exert effort, while the upper time limit was employed to simulate real-life coding conditions, where time constraints are typically present when completing coding tasks and to enable scheduling of sessions.
Figure 2: Overview of page three
On the fourth page, we provided a code snippet and notified the participants that it was generated by an AI code generator. They were also notified that they could modify the code written by themselves on page three with the assistance of the provided AI-generated code. They had the option to use all the provided code, integrate part of the provided code into their original code, or choose not to use the provided code at all. We employed the same timer, code editor, and code execution features on page three.
The fifth page of the survey gathered information on participants’ familiarity with the domain knowledge ($B_{13}$).
On the sixth page, we expressed our gratitude to the participants. We provided a completion code for them and guided them to provide it so that we knew that they had finished the study. After receiving the completion code, we would pay them.
3.2.2 Motivation and Engagement Mechanisms
To motivate participants to focus on doing this survey and put in more effort, we have utilized two methods, including bonus payment and full-screen mode. For the bonus payment mechanism, we informed the participants that their reimbursement depended on the quality of the code they submitted on page three and page four. We told participants that they would only receive 2.6 GBP if the quality of their code was low and that they would receive 1 additional GBP as a bonus award if they submitted high-quality code. The 1 GBP payment increase was significant in this study because it was a 38% increase in basic payment. In the end, we gave every participant both the basic 2.6 GBP reimbursement and the 1 GBP bonus regardless of the quality of their submitted code because it is hard to objectively determine the quality of their submitted code.
There were two purposes for adopting the bonus payment mechanism. The first purpose was to ensure genuine engagement from participants. Performance-based incentives could help increase participants’ performance [29]. The difference in payment could motivate participants to put more effort into their participation. The second purpose was to motivate participants to provide more representative data. Programmers are often required to write high-quality code in their daily work. Therefore, motivating participants to produce high-quality code in our survey was a way to motivate them to approach the task as if they were coding in their actual professional work.
The second motivation mechanism was full-screen mode, which encouraged participants to stay in full-screen mode while doing the survey. We motivated participants to stay in full-screen mode for two reasons. Firstly, staying in full-screen mode helps participants focus on the survey. Secondly, maintaining the survey in full-screen mode helps to prevent participants from seeking external assistance, such as information from the Internet and solutions created by the participants themselves using AI tools. Our goal was for participants to interact solely with the AI-generated solutions provided in our survey to gauge their adoption accurately. External information could influence their adoption of the provided AI-generated code in the survey, which might introduce bias.
We used a dialogue window to guide participants to stay in full-screen mode. After participants signed the consent form, they would see a dialogue window as shown in Fig. 3. This dialogue window informed the participants that they needed to stay in full-screen mode, and their behavior of leaving full-screen mode would be recorded. We also warned them that leaving full-screen mode may lead to no reimbursement, although we still paid them even if they had left full-screen mode. We recorded this data to enable us to discard participant data that may have been tainted by seeking outside assistance. Once participants clicked the OK button, their browsers were switched to full-screen mode. Once they left full-screen mode, this dialogue window shown in Fig. 3 would appear again. This survey design was approved by the ethics review board of our institution and follows the rules of the Prolific data collection platform used in this work.
Figure 3: Full-screen mode dialogue window while not in full-screen mode
3.3 Experimental Objects
The experimental objects contained programming problems, the corresponding predefined test cases, and AI-generated solutions to the programming problems.
3.3.1 Programming Problems
We have chosen five programming problems from LeetCode, a well-known and popular platform where programmers practice programming [30, 31]. Each survey will be randomly assigned one of these five programming problems and the participant needs to solve this programming problem on the third page and the fourth page of the survey. We chose programming problems from LeetCode for the following two reasons. Firstly, LeetCode’s popularity increases the chances of high-quality and correct programming problems. The programming problems from LeetCode have clear introductions and well-crafted content. Moreover, LeetCode programming problems have been proven to be good experimental materials, as they have been used in multiple studies to evaluate AI code generation [25, 32, 33].
To enhance the generalizability of this study, we chose five programming problems with different topics and different difficulty levels. According to the topic tags provided by LeetCode, the topics of these programming problems include map, stack, queue, linked list, and tree. We have chosen two easy programming problems, two medium programming problems, and one hard programming problem based on the difficulty labeled by LeetCode. A brief summary of the five selected programming problems is presented in Table 1.
Table 1: A brief summary of selected programming problems
Name Topics Difficulty Level LeetCode Number*
Ransom Note Map Easy 383
Remove Duplicates from Sorted List Linked List Easy 83
Find the Winner of the Circular Game Queue Medium 1823
Validate Binary Search Tree Tree Medium 98
Longest Valid Parentheses Stack Hard 32
    LeetCode Number is a unique identifier assigned to each problem on the LeetCode platform for easy reference and search.
3.3.2 Predefined Test Cases
Participants were able to execute their code using predefined test cases on both page three and page four of the survey. Each programming problem included 10 pre-defined test cases, all of which were hidden from participants. These test cases were generated by GPT 4 and subsequently reviewed by the authors of this paper to ensure their correctness and scalability.
3.3.3 AI-generated Solutions
We have created two versions of AI-generated solutions for each programming problem: solutions with comments and solutions without comments. Every survey was assigned one of these two versions. On the fourth page of the survey, participants used the assigned AI-generated solutions to assist in modifying the code written by themselves on page three. The provided AI-generated code snippets were generated by ChatGPT 4. To generate the solutions, we copied the problem description, including examples, into ChatGPT 4. Then, we added the prompt ”Please solve this problem in JavaScript” and included the template code from LeetCode, which provides function signatures. We then added a final prompt based on the version: ”Please do not add any comment” for the no-comment version and ”Please add meaningful comments” for the commented version. We confirmed that all the code snippets generated by ChatGPT in this research could be executed correctly and pass all test cases on LeetCode.
3.4 Variables, Measurements, and Analyses
Table 2 demonstrates the variables in this study, including independent variables, control variables, and dependent variables.
Table 2: Outline of variables
Abbreviation Full Name Category
$A_{1}$ Presence of commentsa Independent variable
$B_{1}$ Development expertiseb Independent variable
$B_{11a}$ Perceptions of programming experience Independent variable
$B_{11b}$ Years of overall programming experience Independent variable
$B_{12a}$ Perceptions of programming language proficiency Independent variable
$B_{12b}$ Years of experience in JavaScript development Independent variable
$B_{13a}$ Perceptions of domain knowledge familiarity Independent variable
$B_{13b}$ Prior exposure to domain knowledge Independent variable
$B_{2}$ Genderc Control variable
$B_{3}$ Aged Control variable
$B_{4}$ Attitudes toward AI code generatorse Control variable
$B_{4f}$ Fear of AI code generators Control variable
$B_{4a}$ Attitudinal acceptance of AI code generators Control variable
$Y$ Latent code similarityf Dependent variable
  • a
    The variable ”presence of comments” is manipulated by providing AI-generated code snippets either with comments or without them.
  • b
    The variable ”development expertise” is obtained through PCA analysis on $B_{11a}$, $B_{11b}$, $B_{12a}$, $B_{12b}$, $B_{13a}$, and $B_{13b}$.
  • c
    Data of the variable ”gender” is provided by Prolific.
  • d
    Data of the variable ”age” is provided by Prolific.
  • e
    Attitudes toward AI code generators comprise two aspects, which are fear and attitudinal acceptance.
  • f
    Latent code similarity is computed through PCA analysis on four different code similarity metrics in Table LABEL:tab:similarity-metrics.
3.4.1 Independent Variables
The independent variables in this research were ($A_{1}$) the presence of comments and ($B_{1}$) development expertise. The variable $A_{1}$ was a categorical variable determined by whether the AI-generated code included comments. The value of $A_{1}$ was 1 if comments were present, and the value of $A_{1}$ was 0 if comments were absent.
To capture the complexity of ($B_{1}$) development expertise, the corresponding factor $B_{1}$ has three sub-factors, including ($B_{11}$) programming experience, ($B_{12}$) programming language proficiency, and ($B_{13}$) domain knowledge familiarity. We assessed ($B_{1}$) development expertise by performing PCA (Principal Component Analysis) on $B_{11}$, $B_{12}$, and $B_{13}$. The detailed PCA analysis will be introduced in Section 4.1.1.
To enhance the reliability of the obtained data, we utilized two questions to assess each sub-factor. First, we employed a Likert scale question ($B_{11a}$, $B_{12a}$, $B_{13a}$) for each sub-factor, translating participants’ perceptions into integers from 1 (Not experienced/proficient/familiar at all) to 5 (Extremely experienced/proficient/familiar). For sub-factors $B_{11}$ and $B_{12}$, we used questions $B_{11b}$ and $B_{12b}$ to gauge participants’ years of experience in overall programming and JavaScript development, respectively. The second question used to assess $B_{13}$, question $B_{13b}$, measured participants’ familiarity with specific domain knowledge by determining their level of prior exposure to certain topics. The possible levels were ”I have never heard of x,” ”I have heard of x but never used it,” ”I have occasionally used x,” and ”I have regularly used x” (here, x represents a certain domain knowledge topic).
3.4.2 Control Variables
The control variables are ($B_{2}$) gender, ($B_{3}$) age, and ($B_{4}$) attitudes towards AI Code Generators. We evaluated ($B_{4}$) attitudes towards AI Code Generators by mirroring and modifying the Attitudes Towards Artificial Intelligence (ATAI) scale proposed by Sindermann et al. [34]. Akin to the ATAI scale, we categorized participants’ attitudes toward AI code generators into two domains: fear and attitudinal acceptance. Therefore, we had two sub-factors for $B_{4}$, which are ($B_{4f}$) fear of AI code generators and ($B_{4a}$) attitudinal acceptance of AI code generators. We explored $B_{4}$ through grid Likert scales in Fig. 4, where the mean value of the responses to the first two questions represents $B_{4a}$, and the mean value of the responses to the last three questions represents $B_{4f}$.
Figure 4: Grid Likert-scale question on evaluating participants’ attitudes towards AI code generators [34]
3.4.3 Dependent Variable
To measure the adoption of AI-generated code by programmers, we devised a novel objective metric. Participants were tasked with writing a code snippet with the assistance of an AI-generated solution (page four of the survey). The degree of adoption was then quantified by comparing the similarity between the code submitted by the participants and the AI-generated code provided on page four. This method provides an objective assessment of the code adoption behavior.
To gain insights into the dependent variable from different dimensions, four distinct code similarity metrics were utilized in this study, including line-level code similarity from the participant perspective (LP), line-level code similarity from the AI perspective (LA), token-sequence-level code similarity from the participant perspective (TP), and token-sequence-level code similarity from the AI perspective (TA). These metrics could be categorized by two dimensions: the dimension of entity perspective and the dimension of code similarity level. Table LABEL:tab:similarity-metrics shows the classification and interpretation of these metrics. For data analysis, we created a new latent variable, latent code similarity, by combining LP, LA, TP, and TA through PCA. The detailed PCA analysis is presented in Section 4.1.2. The new latent variable, latent code similarity, was the dependent variable for hypotheses testing.
The dimension of entity perspective indicated the entity from which the similarity metric was assessed, suggesting which entity was used as the denominator when calculating similarity. The possible values for the entity perspective dimension were participant and AI. If the entity perspective of a code similarity metric is ”participant,” then the metric represents the percentage of the participant’s code that is similar to the AI-generated code, highlighting to what extent the participant’s coding approach has been influenced by the AI-generated solution. If the entity perspective of a code similarity metric is ”AI,” then the metric represents the percentage of the AI-generated code similar to the participant’s code, revealing the percentage of AI-generated code used by the programmer.
The dimension of code similarity level showed the level at which the comparison was conducted. The possible values for the dimension of code similarity level were line level and token sequence level. The line-level code similarity metrics calculated the ratio of identical lines to total lines. Only lines that were exactly the same in both the AI-generated code and the participants’ submitted code would be counted as a match (differences in whitespace and comments were ignored). Token-sequence-level code similarity metrics compute similarity by comparing the token sequences of two code snippets, a process carried out using Jplag [35]. The token is a basic element of the code abstraction [35, 36]. A token sequence represents some fractions of the code snippet without necessarily representing a line of code. Token sequence abstracts the code, providing a mapping to the code structure without being influenced by different identifier names and comments [36]. Some code similarity-checking tools were developed to detect plagiarism in students’ programming assignments at a token-sequence level [37], including MOSS [36] and JPlag [35]. We used JPlag [35] to calculate the code similarity at the token-sequence level  [^1] We utilized the JPlag Java package from https://github.com/jplag/JPlag to calculate the token-sequence-level similarity. Line-level code similarity investigated participants’ direct copy-and-paste behavior. In contrast, the token-sequence-level code similarity focused on the abstract structure of the code.
Table 3: Classification of code similarity metrics
Abbreviation Entity Perspective
Dimension
Code Similarity Level Dimension Interpretation
LP Participant Line level Percentage of the participant’s code that are similar to the AI-generated code at the line level.
LA AI Line level Percentage of the AI-generated code that is similar to the participant’s code at the line level.
TP Participant Token sequence level Percentage of the participant’s code that are similar to the AI-generated code at the token sequence level.
TA AI Token sequence level Percentage of the AI-generated code that are similar to the participant’s code at the token sequence level.
Table 3: Classification of code similarity metrics (continued)
3.5 Participants
Participants in this study were recruited from Prolific [38], an online crowd-sourcing academic data collection platform used extensively in multiple research fields [38], including software engineering [8, 39, 40]. This study strictly followed the applicable guidelines for ethical conduct in human research and was approved by the ethics review board of our institution.
3.5.1 Pre-Screening Mechanism
Participants were pre-screened to ensure they met the study’s eligibility criteria. We utilized pre-screeners provided by Prolific to select participants with programming experience in JavaScript to ensure they were capable of doing the programming tasks in the survey. To minimize misunderstanding of the English instructions in our survey, we further pre-screened participants to ensure their primary language was English. The last pre-screener excluded participants with approval rates below 95%. The approval rate indicates the percentage of submissions approved by researchers. A low approval rate could indicate a participant’s history of negative behaviors, such as submitting low-effort responses [41]. To mitigate low-effort responses in our survey, we required participants to have at least a 95% approval rate.
3.5.2 Post-Survey Screening Mechanism
We employed two post-survey screening approaches to filter out invalid responses. The first approach was excluding responses that might have sought external assistance. If participants left full-screen mode more than twice or for longer than two seconds, then they were deemed as having sought external assistance. As illustrated in Section 3.2.2, using external resources could introduce bias to participants’ adoption of the provided AI-generated code. Therefore, these responses were viewed as invalid responses and excluded from the data analysis.
The second post-survey screening approach was to exclude low-effort responses. A low-effort response has a damaging impact on survey-based research as it introduces untrustworthy data [42]. To examine the level of effort of a response, we manually reviewed the code participants submitted on page 4. We examined whether the complexity of the submitted code was significantly low. For instance, when required to validate whether a tree is a valid binary search tree, a participant submitted a solution that directly returned true. This participant returned a hard-coded value rather than trying to solve the programming problem. An overly simplistic code that did not address the programming problem suggested low effort.
In this study, we collected responses from 251 participants in total. After employing the post-survey screening mechanism, we obtained 173 valid responses. The demographics of these valid respondents are showcased in Table LABEL:tab:demographics, which highlights the diversity of the participant sample.
Table 4: Demographic Summary of Survey Participants
Characteristic Description
Gender Male (N = 119), Female (N = 54), Others (N = 0)
Age Minimum: 18, Maximum: 68, Mean: 27, SD: 7.748; Age distribution: $<$23 (25%), 23–32 (50%), $>$32 (25%)
Country of Residence South Africa (N = 52), United Kingdom (N = 18), Canada (N = 16), Portugal (N = 16), Poland (N = 10)
Country of Birth South Africa (N = 46), United Kingdom (N = 14), Portugal (N = 14), Poland (N = 10), Canada (N = 8)
Employment Status Employed full-time (N = 91), Employed part-time (N = 21), Unemployed (N = 31), Other (N = 30)
Occupation Professional specializing in IT (N = 80, 46%), Student specializing in IT-related fields (N = 53, 30.5%), Others (N = 41, 23.6%)
Years of Programming Experience Mean = 4.13 years, Standard deviation = 4.365; 25th percentile: 2 years, 50th percentile: 3 years, 75th percentile: 5 years
Years of JavaScript Experience Mean = 4.13 years, Standard deviation = 4.365; 25th percentile: 1 year, 50th percentile: 1.5 years, 75th percentile: 3 years
Table 4: Demographic Summary of Survey Participants (continued)
4 Results
4.1 Preliminary Analysis
Before testing the hypotheses, we first calculated the latent variables ($B_{1}$) development expertise and ($Y$) latent code similarity.
4.1.1 Development Expertise
We created the latent variable development expertise ($A_{3}$) in two steps. Firstly, we used Cronbach’s alpha [43] to test whether all the questions used for assessing programmer factors were measuring the same underlying construct. Secondly, we used principal component analysis (PCA) [44] to create the latent variable ($B_{1}$) development expertise with observed variables $B_{11a}$, $B_{11b}$, $B_{12a}$, $B_{12b}$, $B_{13a}$, and $B_{13b}$.
Cronbach’s alpha is a measure of reliability that determines the internal consistency of multiple items designed to measure the same construct [43]. The value of Cronbach’s alpha of $B_{11a}$, $B_{11b}$, $B_{12a}$, $B_{12b}$, $B_{13a}$, and $B_{13b}$ was 0.833, higher than the cut-off value 0.8 [43]. This result indicated a strong internal consistency among all the scales of observed programmer factors.
We then performed PCA on $B_{11a}$, $B_{11b}$, $B_{12a}$, $B_{12b}$, $B_{13a}$, and $B_{13b}$ to compute ($A_{3}$) development expertise. We tested the linearity assumption [45] and the sampling adequacy assumption [46] to ensure that the collected data were suitable for PCA. We then used the scree plot criterion [45] to choose latent factors that represent the observed variables. The scree plot criterion [45] suggested that only the first factor should be chosen. Keeping the first component explained 54.729% of the variance, which was acceptable [45]. The component matrix is shown in Table 5, indicating that the chosen factor could cover all observed variables. The interpretation of the data was consistent with the concepts of development expertise, since it involves positive coefficients for overall programming experience, programming experience in particular programming languages, and task-specified knowledge familiarity.
Table 5: Component matrix for development expertise
Observed Variable $B_{11\text{a}}$ $B_{11\text{b}}$ $B_{12\text{a}}$ $B_{12\text{b}}$ $B_{13\text{a}}$ $B_{13\text{b}}$
Coefficient 0.751 0.813 0.748 0.758 0.722 0.635
4.1.2 Latent Code Similarity
To simplify the data analysis and see the big picture of code similarity, we used PCA to generate the independent variable latent code similarity inferred from the four code similarity metrics listed in Table LABEL:tab:similarity-metrics. The procedure of computing latent code similarity was the same as computing development expertise. One latent factor was selected to represent the TP, TA, LP, and LA. The retained factor explained 82.550% of the total variances. As shown in Table 6, the retained latent variable could capture all the features of the four different code similarity metrics.
Table 6: Component matrix for latent code similarity
Code Similarity Metric TP TA LP LA
Coefficient 0.915 0.905 0.896 0.918
We also observed the difference between participants’ code written on page three and page four. There were seven participants who did not modify their code after being provided with the AI-generated code. This result indicated that the majority of participants (95.95%) were influenced by the AI-generated solution.
4.2 Hypotheses Testing
We employed linear regression to examine the influence of ($A_{1}$) presence of comments and ($B_{1}$) development expertise on (Y) latent code similarity, using a data analysis approach consistent with previous studies [23, 47]. We hypothesize that the level of development expertise exerts a moderation effect [48] on the influence of comments on programmers’ adoption of AI-generated code because more experienced programmers may rely less on comments to understand and integrate the code, while less experienced programmers may find comments crucial for comprehension and adoption. To explore the potential moderation effect, we added an interaction term into the linear regression models. The interaction term was calculated as the product of ($A_{1}$) presence of comments and ($B_{1}$) development expertise, denoted as $A_{1}*B_{1}$.
We firstly included ($A_{1}$) presence of comments, ($B_{1}$) development expertise and the interaction term $A_{1}*B_{1}$ into the linear regression. The overall model significantly predicted ($Y$) latent code similarity, $F(3,173)=2.724,p=0.46$. The coefficient table for the factors $A_{1}$, $B_{1}$, and $A_{1}*B_{1}$ is shown in Table 7. The results demonstrated that ($A_{1}$) the presence of comments could significantly increase programmers’ adoption of AI-generated code ($\beta=0.336,p=0.027$), supporting hypothesis $H_{1}$. We anecdotally observed that comments were copied in the majority of cases and that participants also tended to copy function signatures. A systematic analysis of these behaviors is part of our future work.
As showcased in Table 7, neither ($B_{1}$) development expertise nor the interaction term between ($B_{1}$) development expertise and ($A_{1}$) presence of comments had a significant influence on code adoption ($p>0.05$). This result indicated that the level of development expertise did not moderate the influence of ($A_{1}$) presence of comments. Programmers tended to adopt more AI-generated code when comments were present, regardless of their development expertise. The insignificance of $B_{1}$ and $A_{1}*B_{1}$ suggested that ($B_{1}$) development expertise did not diminish the positive effect of ($A_{1}$) presence of comments on programmers’ adoption of AI-generated code. Therefore, we reject hypothesis $H_{2}$.
Table 7: Coefficients and Significance of Key Factors
Variable Coefficient Significance
($A_{1}$) presence of comments 0.336 0.027
($B_{1}$) development expertise -0.113 0.253
($A_{1}*B_{1}$) interaction term -0.015 0.922
To further confirm and validate the results in Table 7, we added all control variables, including ($B_{2}$) gender, ($B_{3}$) age, ($B_{4f}$) fear of AI code generators, and ($B_{4a}$) attitudinal acceptance of AI code generators, into our initial linear regression model. The coefficients and significance of all variables included in the second model are presented in Table 8. Table 8 indicated that ($A_{1}$) presence of comments was a significant predictor of (Y) latent code similarity, while ($B_{1}$) development expertise and the interaction term $A_{1}*B_{1}$ remained insignificant. The second linear regression model confirmed the results of the initial regression model: even when accounting for all control variables, the presence of comments increases programmers’ adoption of AI-generated code regardless of programmers’ development expertise. Table 8 also showed that all control variables were insignificant, including participants’ attitudes towards AI code generators. The discrepancy between self-reported attitudes and actual behavior demonstrates the limitation of overly relying on self-reported intention of adopting AI coding tools, further highlighting the importance of adopting objective measurements of programmers’ code adoption behavior.
Table 8: Coefficients and Significance of Key Factors and Control Variables
Variable Coefficient Significance
($A_{1}$) presence of comments 0.354 0.022
($B_{1}$) development expertise -0.138 0.176
($A_{1}*B_{1}$) interaction term -0.021 0.889
($B_{2}$) gender 0.117 0.490
($B_{3}$) age 0.009 0.401
($B_{4f}$) fear of AI code generators 0.003 0.921
($B_{4a}$) attitudinal acceptance of AI code generators -0.047 0.280
5 Discussion and Implications
Our research has shown that comments, an essential element in facilitating code comprehension, continue to play a significant role in influencing programmers’ adoption of AI-generated code. The presence of comments significantly encourages programmers to adopt AI-generated code, regardless of their development expertise. Both novice and expert programmers tended to adopt more AI-generated code when it included comments.
For novice programmers, the rationale for the influence of comments could be explained in the following way: the presence of comments enhances code comprehensibility, and better code comprehensibility results in a stronger intention to adopt AI-generated code.
Previous studies have demonstrated the profound influence of comments on code comprehension [21, 47]. Kolodziej discovered that code snippets with comments significantly decreased the time needed for code comprehension and improved the correctness rates in code comprehension tests [47]. Börstler and Paech evaluated the influence of comment quality on code comprehension, and found that even comments of poor quality can lead to a higher perceived code comprehension compared to code snippets without comments [21].
The comprehensibility of AI-generated code plays an important role in participants’ adoption of AI-generated code. Previous studies have shown that many users perceived code understandability as a limitation of AI coding tools [5, 49, 10]. GitHub Copilot user studies discovered that users found it difficult to understand the code generated by Copilot [5, 49], and they did not use Copilot to perform programming tasks due to the difficulty in understanding the Copilot-generated code [5]. Liang et al. conducted a large-scale survey to evaluate the factors that influence the adoption and rejection of AI coding assistants [10]. Their study indicated that 16% of the participants rated ”I don’t understand the code written by code generation tools” as ”Very important” or ”Important” when asked about reasons for not using AI coding tools [10]. The limitation of unsatisfying code comprehensibility was one of the reasons programmers did not use AI coding tools, and poorly written documentation or the absence of comments could be the reason for unsatisfying code understandability.
Our study revealed that the presence of comments can also increase the adoption of AI-generated code for expert programmers. Previous studies have discovered that expert programmers were more capable of understanding source code [23, 24]. It is reasonable to extrapolate that expert programmers were more capable of understanding the AI-generated code without comments. However, they were still more likely to adopt AI-generated code that was commented. The rationale for the influence of comments on expert programmers’ code adoption might not be fully explained by the impact of comments on code comprehensibility. One possible reason why expert programmers prefer commented code is that commented code often indicates higher code quality, and expert programmers may have more stringent requirements for code quality. Code comments could be leveraged to improve the readability, maintainability, and reliability of software [50]. Expert programmers are expected to produce high-quality code that exhibits several desirable characteristics, including enhanced readability and strong maintainability [22]. Commented code was more capable of meeting the requirements of expert programmers.
To encourage more programmers to utilize AI coding tools and enhance the adoption rates of AI-generated code, the developer of AI coding tools should consider ensuring the presence of comments in generated code snippets and increasing the quality of accompanying documentation.
6 Limitations and Threats to Validity
Construct Validity. The measurement of development expertise might not accurately reflect the true level of participants’ development expertise. Development expertise was quantified as a linear combination of ($A_{31}$) programming experience, ($A_{32}$) programming language proficiency, and ($A_{33}$) domain knowledge familiarity. Although we incorporated ostensibly objective measures, such as the number of years participants had focused primarily on programming in their work or studies, these metrics still relied on self-reporting, which can be inherently subjective and potentially unreliable. Baltes and Diehl investigated the validity of self-assessed expertise ratings and years of programming experience among various groups of programmers [22]. They concluded that neither the self-evaluated experience ratings nor the years of experience could reliably measure programmers’ actual expertise [22].
Measuring AI-generated code adoption through similarity metrics can be objective but may miss some aspects of adoption. While our metrics capture the adoption of code structures and exact lines of code, they overlook the adoption of stylistic aspects, such as identifier naming styles, variable naming conventions, and overall code structuring preferences. Specifically, TP and TA ignore identifier name changes, and LP and LA assess only exact line matches. Thus, if a participant adopts naming styles of the AI-generated code, like the camel case naming style, or follows function decomposition patterns present in the AI-generated code, our metrics will not capture this. Additionally, our metrics also fail to capture the situation where programmers might unconsciously adopt AI-generated code structures, for instance, having a preference for arrow functions instead of traditional function expressions when working with JavaScript. Future research could incorporate style similarity analysis and investigate whether prolonged exposure to AI-generated code affects programmers’ personal coding styles over time.
External Validity. This study faces several threats to external validity. Firstly, the programming language was limited to JavaScript. We chose JavaScript for practical reasons, as using multiple languages would have required significantly more participants and resources. Previous studies have shown that the quality of AI-generated code varies across different languages [33]. Limiting the study to only one programming language may not accurately capture the overall characteristics of AI-generated code in other languages. Future research should explore how AI-generated comments influence adoption across multiple programming languages, considering that different languages have varying conventions for code readability, documentation practices, and idiomatic structures.
Secondly, the scope of the programming tasks was narrow and restricted to data structure and algorithm problems from LeetCode. We selected LeetCode for feasibility, as it is widely used in related studies [25, 32, 33, 30, 31] and provides a familiar, controlled platform for experimentation. Although these problems may not fully represent real-world tasks, they allowed us to cover a range of difficulty levels and topics. This restriction may not accurately represent real-world software engineering contexts, such as web development, system design, and database integration. Additionally, the difficulty of the tasks could influence programmers’ adoption of AI-generated code. While we included problems of varying difficulty, we did not analyze whether programmers were more likely to adopt AI-generated code for easier or harder problems. It is possible that programmers struggling with complex problems may rely more heavily on AI-generated solutions, while those solving simpler problems may feel more confident in their own implementations. Future research should systematically explore how task complexity affects AI-generated code adoption.
Thirdly, the complexity of the programming tasks was constrained, with a limited number of possible solutions and short solution lengths, which may not capture the open-ended nature of real-life software development tasks. In professional settings, developers often work on projects with multiple valid approaches, varying levels of abstraction, and domain-specific constraints. Further research could explore AI-generated code adoption in larger-scale software projects where contextual factors such as maintainability, codebase integration, and long-term readability play a role.
Fourthly, the study used error-free AI-generated code, which did not reflect real-world scenarios where AI-generated code can contain errors and vulnerabilities [51, 52]. In real-world applications, developers frequently need to evaluate AI-generated code for correctness and security before adopting it. Additionally, AI-generated code might contain subtle inefficiencies or biases in implementation that could impact long-term maintainability. Future studies could investigate how the presence of AI-generated errors influences adoption behavior, particularly whether programmers with higher expertise are more selective in integrating AI-generated solutions.
Lastly, participants from Prolific are not necessarily representative of programmers, although we tried to mitigate this issue through pre-screening mechanisms. While our participants included those with full-time employment in IT and computer science students, the recruitment method may have introduced biases compared to recruiting professional software engineers working in the industry. Future research could replicate this study with industry practitioners to assess whether similar adoption patterns emerge in workplace settings.
Future studies could enhance the generalizability of the results by expanding the scope and complexity of programming tasks, evaluating the use of erroneous AI-generated code, and recruiting a more diverse group of programmers through professional networks.
Internal Validity. To maintain procedural control, we implemented methods to encourage participants to adhere to the designed procedures. One potential threat is related to the difficulty in preventing participants from copying solutions generated by external AI. To address this, participants were initially asked to solve each problem independently to motivate active engagement and filter out low-effort responses. This self-exclusion approach helped improve data quality by filtering out less meaningful responses.
While acknowledging that external factors could have impacted the code produced, our primary focus remained on adoption behavior, which we believed to be relatively unaffected by these external influences. Even if participants used external AI tools to verify their work, they were still less likely to adopt code without comments, which stands as our central finding, underscoring the influence of comments on code adoption. However, external AI assistance could introduce biases, particularly if participants cross-referenced different AI-generated solutions before submitting their final code. Moreover, different AI tools may produce stylistically distinct solutions, and the AI model used in this study may exhibit implicit biases in generating code. Future studies should examine how different AI code generators influence programmers’ adoption behaviors and whether developers prefer AI-generated solutions from specific models based on readability or correctness.
Additionally, real-world software development is often constrained by deadlines, code review processes, and collaboration with other developers, all of which could influence code adoption. Our controlled environment did not fully replicate these constraints, and future research could examine how time pressure and team collaboration affect programmers’ willingness to adopt AI-generated code. Furthermore, it remains unclear whether repeated exposure to AI-generated solutions could lead to over-reliance on AI suggestions or whether programmers develop strategies to critically assess AI-generated code over time. Future studies could investigate the long-term effects of AI-assisted coding on software development practices.
7 Conclusion and Contributions
Our survey-based study explored how the presence of comments and development expertise influence programmers’ adoption of AI-generated JavaScript code. Participants reported their expertise, attempted to solve a programming problem independently, and then received AI-generated solutions with or without comments. Adoption was objectively measured through code similarity between participants’ submissions and the AI-generated solutions. After data collection, we applied principal component analysis and linear regression to analyze the effects.
Our study has made several important contributions. Firstly, we examined unexplored factors that could potentially affect programmers’ adoption of AI-generated code. We found that programmers adopted more AI-generated JavaScript code when comments were present, regardless of development expertise. For industry practitioners, this finding has the potential to motivate developers of AI coding systems to place greater emphasis on annotating AI-generated code. For instance, future training in AI programming systems could incorporate more code with high-quality comments into the training set. This can help create more user-friendly AI coding tools, further boosting the productivity of AI coding assistants. For academia, this finding allows researchers to have a better understanding of the AI-human collaboration paradigm. Future studies could further expand this finding by evaluating what types of comments can help enhance programmers’ adoption of AI-generated codes.
Secondly, we proposed novel objective measurements for programmers’ adoption of AI-generated code, which are valuable complements and alternatives to traditional self-reported data. Rather than directly inquiring of participants about the degree to which they adopted AI-generated code, we employed a more objective approach. We measured the degree of code adoption by calculating the similarity between the participants’ submitted code and the AI-generated code. We evaluated code similarity from two distinct dimensions: the entity perspective and the code similarity level. This new approach could provide more objective and comprehensive insights about programmers’ adoption of AI coding tools. We have also provided the replication package [^2] The data preprocessing scripts, raw data, and the data analysis results can be found at: https://github.com/ArthurLCW/CodeResponseAnalysisScript to help future researchers utilize our measurements of code adoption.
Thirdly, we developed a custom survey application that facilitated coding tasks [^3] To access the survey app, please visit: https://survey.changwen-software-engineering.xyz; Source code of the survey app can be found at: https://github.com/ArthurLCW/CodeResponseSurveyWebsite. This survey application incorporated a code editor with advanced features, including syntax highlight and auto-completion, providing a coding experience similar to using VSCode [27]. This allowed participants to complete coding tasks in a familiar environment. Future survey-based research can adapt and extend this application to suit their specific requirements, especially if the surveys include coding components.